OpenStudy (anonymous):

Use the Euclidean algorithm to express the greatest common divisor of each of these pairs of integers as a linear combination of these integers. c) 35, 78

myininaya (myininaya):

I could give you an example and you could ask me if you don't understand something and I will explain it the best I can.

myininaya (myininaya):

Example: Let's first find the gcd of 18 and 132. so 132 is bigger so we will start with it. 132=18( 7)+6 Note 132/18=7+6/18 18=6(3)+0 <stop when remainder is 0 so the gcd is the previous remainder so the gcd of 18 and 132 is 6. also let's write 6 as a linear combination of 18 and 132. so we had above that 132=18(7)+6 solve for 6 132-18(7)=6 done! now some examples could be longer i could do another one but let's see if you can at least find the gcd using eculid

myininaya (myininaya):

@academicpanda which of your numbers is bigger?

OpenStudy (anonymous):

I have this: 1=3-2 1=3-(8-2*3) 1=3-8+2*3 1=-8+2*3+3 1=3*3-8,as 2*3 is the same as 3+3 and 3+3+3 is the same as 3*3 1=3*(35-4*8)-8 1=3*35-12*8-8 1=3*35-13*8 1=3*35-13*(78-2*35) 1=3*35-13*78+26*35 1=29*35-13*78

myininaya (myininaya):

well yeah that is right 29*35-13*78 is totally 1 so good job

myininaya (myininaya):

I guess you had no problem with the gcd part?

myininaya (myininaya):

I assume that since you were able to go backwards through that process to write 1 as a linear combination of 35 and 78.

OpenStudy (anonymous):

yeah i wasn't sure i was using Euclidean's right lol

myininaya (myininaya):

Well now you know. Sometimes I seen a lot of people get lost in finding the linear combination thing when you have a lot of steps in finding the gcd

myininaya (myininaya):

it is just basically going backwards through the gcd part rewriting the remainder as the "subject"

myininaya (myininaya):

With a lot of substituting other numbers for other numbers

myininaya (myininaya):

i mean they are technically the same number just written different lol

OpenStudy (anonymous):

cool thanks!